OpenStudy (anonymous):

Help if you can!!! I will MEDAL best explanation/answer!!!! Solve for x. -ax + 2b > 8

OpenStudy (anonymous):

oh no not another one!!

OpenStudy (anonymous):

haha yea I guess so! :/ XD

OpenStudy (anonymous):

I just really need help on this one, like a detailed explanation on how to solve for x..

OpenStudy (anonymous):

you cannot, i repeat CANNOT solve an inequality with variable coefficients i will tell you what they want you to say, but only after if you tell me where the question comes from whoever wrote it is either extremely careless or doesn't know any math

OpenStudy (anonymous):

Umm, it's from one of my math quizzes for school.. I didn't write it, the school did lol

OpenStudy (anonymous):

I don't know why they give it to us if it's wrong but I don't have a say in what problems they give us.. :/

OpenStudy (anonymous):

i know you didn't i will answer you when you tell me exactly where the question comes from on line school? in class school, where exactly

OpenStudy (anonymous):

from an online school called MEWA... it's a Algebra review quiz and I am stuck on how to solve it

OpenStudy (anonymous):

Is that detailed explanation enough?? I don't know what else to say about where it came from.. :/

OpenStudy (anonymous):

you got an item number for this question?

OpenStudy (anonymous):

what is MEWA stand for?

OpenStudy (anonymous):

Metro East Web Academy

OpenStudy (anonymous):

we have answer choices but it is in image form and I don't know how to post it

OpenStudy (anonymous):

ok i will write to them now here is what they want you to do

OpenStudy (anonymous):

hahaha okayy but don't mention my name! haha

OpenStudy (anonymous):

\[-ax+2b>8\]first subtract \(2b\) from both sides , get \[-ax=8-2b\]

OpenStudy (anonymous):

okay, that's what I was thinking I had to do... do you have to divide A from both sides?

OpenStudy (anonymous):

then they want you to do this ( and this is the mistake) they want you to divide by \(-a\) and say "since \(-a\) is negative, i should switch the inequality" and write \[x<\frac{8-2b}{-a}\] or \[x<\frac{2b-8}{a}\]

OpenStudy (anonymous):

that is what they want from you hope the steps are clear

OpenStudy (anonymous):

Yea that's what they taught us to do, it's a specific rule of theirs..

OpenStudy (anonymous):

ok good luck

OpenStudy (anonymous):

Yea they are! and thank you, that was pretty easy :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thanks!!